An online medical advice company just completed an IPO with an investment bank on a firm-commitment basis. The firm issued five million shares of common stock, and the underwriting fees were $2.00 per share. The offering price was $25.00 per share.

What were the total proceeds from the common-stock sale?
Total proceeds $

How much money did the company receive?
Net proceeds to firm $

How much money did the investment bank receive in fees?
Underwriting spread $
Answer #1

a. Total proceeds from issue = $25.00 × 5,000,000 = $125,000,000

b. Net proceeds to firm = ($25.00 – $2.00) × 5,000,000 = $115,000,000

c. Underwriting spread = $2.00 × 5,000,000 = $10,000,000